You invested $7,500 in a savings account that pays interest of 3.6% compounded monthly. What is the value of your investment after 17 months?

a. $7,365.70   b. $13,682.87    c. $7,891.82    d. $7,541.45    e. None of these

 

12. A non-interest-bearing promissory note has a face value of $950. Find the proceeds of this note if it is discounted 3½ years before its maturity date at 8% compounded quarterly.

a.  $719.98      b.  $725.67     c. $771.65     d. $657.22         e. None of these

 


1
Expert's answer
2021-02-25T05:06:26-0500

Solution.

1) The formula for calculating the value of investment:



"B=A(1+\frac{p}{100})^n,"


where "B" is the future value;

"A" - current value;

"p" - interest rate for the settlement period (day, month, year, ...);

"n" is the number of settlement periods.

Сalculate the amount after 17 months:

"B=7,500(1+\frac{3.6}{100})^{\frac{17}{12}}=\newline
7,500*1.036^{\frac{17}{12}}=7,885.35"$

Answer. e None of these

2) The formula for calculating the value of the note:



"P=S(1-\frac{d}{m})^{m\cdot n},"

where "P" is the value of the note;

"S" - face value;

"d" - interest rate for the settlement period (day, month, year, ...);

"m" is number of charges per year.

"P=950(1-\frac{0.08}{4})^{4\cdot 3.5}=715.96" $

Discount of this note "S-P=950-715.96=234.04" $

Answer. e None of these